# Christian Ghiglino    Alastair Langtry

### v4

@@ 5 The effects of changes in inequality
Having examined changes in the level of income, we now turn to changes in the dispersion of income; that is, changes in inequality. When focusing on a setting with distinct communities (as we did in Section 4 and continue to do so here), there are two dimensions to income inequality: inequality between different communities, and inequality within a given community. In the interest of presenting clean results, we will examine these dimensions one by one. Each dimension yields complementary insights.
Looking at between-community inequality (where we abstract away from within-community inequality) highlights the role of network density. There, communities with denser networks are more sensitive to changes in their income, So redistributing incomes from one community to another increases total consumption when the receiving community has a denser network. And this then pushes communities not party to the redistribution to cut their consumption. This is driven by the substitutability between own consumption and identity group status discussed earlier.
Looking at within-community inequality (where we impose some additional structure on the network in the interest of tractability) highlights how inequality affects consumption. An increase in inequality reduces total consumption, but raises consumption in the communities where income did not change. Again, this is driven by the substitutability between the sources of status.
5.1 Between community inequality
In order to focus on the between-community dimension of inequality, we will assume here that all agents within a given community have the same income. Formally, that $w_{i}=w_{n}$ for all $i\in n$. With this in place, we can now define the ‘density’ of a community. This will be a key metric when studying the effects of changing between-community inequality.
@@ 5.1 Between community inequality
$D_{n}=\sum_{j,k\in n}\left[\sum_{t=0}^{\infty}\widehat{G}^{t}\right]_{jk}$
Readers familiar with Bonacich centrality will notice that our definition of ‘density’ is the sum of centralities of all agents in the community (using the ‘standard’ definition of Bonacich centrality that only involves the network, rather than our variant that also includes income). It is not network density as conventionally defined. [^19] See Jackson (2008, S.2.1.8) for the conventional definition. Nevertheless, it captures an important aspect of how dense connections in a community are on aggregate.
We focus on a particularly simple form of redistribution – a transfer from one community to another, keeping average income in society constant. It reduces income inequality if the community receiving the transfer is poorer than the one giving the transfer, and similarly increases income inequality if the receiving community is richer. We will not consider the ambiguous case where the transfer changes which community is richer. It is clear that much richer changes to the income distribution can be built from a series of these changes.
@@
Consider two communities, $n$ and $n^{\prime}$, both with identity $\theta$. And consider a transfer of income from $n$ to $n^{\prime}$ such that the ranking of the two communities’ incomes does not change.
@@
    Consumption falls in community $n$ and rises in community $n^{\prime}$,
@@
    Total consumption rises if and only if $D_{n^{\prime}}>D_{n}$,
  • (iii)
    Consumption on all other islands rises if and only if $D_{n^{\prime}}<D_{n}$.
The first part of this result is unsurprising. When everyone in a community has higher income, they consume more. And they consume less when they have lower income. But interestingly, the impact of this transfer on overall consumption does not depend on whether it increases or decreases inequality. Rather, it depends on the relative densities of the networks within the two communities that are party to the transfer. If the community receiving the transfer has a denser network than the community giving the transfer, then overall consumption rises. This is because centrality is higher in the receiving community. Therefore consumption there is more sensitive to a change in income than in the giving community. So the increase in consumption in the receiving community more than offsets the fall in the giving community.
This raises total consumption across the two communities – in turn raising the status of the identity group. As before, other (unaffected) communities respond to an increase in the status of their identity groups by cutting their own consumption – due to the substitutability between the two sources of status. But note that this cut in consumption by the other (unaffected) communities cannot outweigh the initial rise in consumption across the two affected communities.
5.2 Within community inequality
In order to focus on the within-community dimension of inequality, we impose a regularity condition on the network structure. Specifically, we assume that each agent has the same number of neighbours (in the weighted sense) with those of each different income level. For example, if there are two different levels of income, $w_{1}$ and $w_{2}$, then each agent has $L$ neighbours with income $w_{1}$, and $L$ neighbours with income $w_{2}$. [^20] Note that if every agent in the network has a unique income level, then this assumption requires that the network is complete. But we believe that there being a small number of different levels of income is a reasonable first-order approximation of reality. Note that this is stronger than assuming the network is regular (which would only require that each agent has $L$ links in total, regardless of the neighbours’ incomes). Formally,
Assumption 3.
Let community $n$ have $Z$ different levels of income, $\{w_{1},...,w_{Z}\}$. Then $\sum_{k}\widehat{G}_{jk}\cdot\mathbf{1}\{w_{k}=w_{m}\}=L$ for all $z$ and for all $j\in n$.
Intuitively, 3 ensures that how strongly $j$ is connected to others is unrelated to how much income those other people have. Colloquially, she is, on average, connected to “Mr and Mrs Average”. This retains the network effects, but “smooths them out” so that they apply equally throughout the network.
Again, we focus on a the simplest possible form of redistribution. Here, that is transfer of income from agents with income $w_{z}$ to agents with income $w_{z^{\prime}}$. This is the same as our transfer in the previous subsection, except that now the transfer happens between sets of agents inside the same community – rather than between two different communities. It reduces income inequality if the agents receiving the transfer are poorer than the agents giving the transfer (and similarly increasing inequality if the receiving agents are richer). As before, we will not consider the ambiguous case where the transfer changes which set of agents is richer.

### v7

@@ 5 The effects of changes in inequality
We now turn to our fifth stylised fact – a puzzling empirical regularity documented by Charles et al. (2009) regarding the effects of income inequality. Specifically, they find that higher income inequality among Whites in a given U.S. state correlates with lower consumption of visible (status) goods by Whites, even when controlling for individual income. But for Blacks, the same rise in inequality leads to higher consumption of visible (status) goods, again holding individual income constant.
In this section, we present an explanation for this stylised fact. There are three steps. First, we present a formal result linking the qualitative impact of changes in inequality to the density of social links in communities. Second, we argue that there is convincing evidence that density is linked to income and has a U-shaped relationship with income. Finally, we show that coupling this U-shaped relationship with our formal result can explain the stylised fact discovered by Charles et al. (2009).
Inequality and the role of density.
To help clarify things, we assume that all agents in a given community $n$ share the same income, expressed formally as $w_{i}=w_{n}$ for all $i\in n$. With this assumption in place, we define the density of a community, which will be a crucial metric for studying the effects of changes in between-community inequality.
@@
$D_{n}=\sum_{j,k\in n}\left[(I-\widehat{G})^{-1}\right]_{jk}$
Readers familiar with Bonacich centrality will recognize that our definition of density sums the centralities of all agents in the community – using the standard definition of Bonacich centrality that focuses solely on network connections, rather than our variant that incorporates income. While this definition diverges from the conventional notion of network density, it captures an important aspect of the overall connectivity within a community. [^15] Conventionally, density is simply the sum of all links (normalised by the number of agents). Here, this would be: $\frac{1}{|n|}\sum_{j,k\in n}\widehat{G}_{jk}$. Both this standard measure and our measure are strictly increasing in the weight of each link.
We focus on a natural form of redistribution – a transfer of income from one community to another while maintaining constant average income across society. Such a transfer decreases income inequality if the receiving community is poorer than the donor community and increases inequality if the recipient is wealthier.
@@
Consider two communities, $n$ and $n^{\prime}$, both sharing the same identity $\theta$. When income is transferred from $n$ to $n^{\prime}$ without changing their income rankings, we observe:
@@
    Consumption decreases in community $n$ and increases in community $n^{\prime}$.
@@
    Consumption in all other communities with identity $\theta$ increases if and only if $D_{n^{\prime}}<D_{n}$.
The first part of this proposition is intuitive: when income rises, consumption increases, and when it falls, consumption decreases. However, the effect of this transfer on the sum of status good consumption in communities $n$ and $n^{\prime}$ depends on their relative densities. If the receiving community ($n^{\prime}$) has a denser network, the increase in consumption there outweighs the fall in the donor community ($n$). This is because a denser network makes the receiving community’s consumption more sensitive to a change in income. This enhances the status of the identity group.
The other communities with identity $\theta$ – those not party to the transfer – respond to this enhanced group status by reducing their consumption. [^16] Importantly, this reduction cannot fully offset the initial rise in consumption among the two affected communities. So we also have that total consumption (across all communities) rises if and only if $D_{n^{\prime}}>D_{n}$. In other words, with higher status from their group, agents feel less need to consume status goods themselves. This is due to the imperfect substitutability between own status good consumption and group status.
Notice that this depends only on the relative incomes of the two communities who are party to the transfer – not on where in the overall income distribution these two communities sit. So it applies equally to changes concentrated at the top of the income distribution (e.g. the very rich pulling away from the merely quite rich) as it does to changes between the rich and the poor.
The link between density and income.
The connection between this proposition and the observed stylised facts hinges on an assumption that density (a la Definition 3) follows a U-shaped pattern relative to income. That is, both low- and high-income communities have the highest densities and so experience the strongest pressures to conform. We argue that there is convincing evidence in support of this pattern.
First, Bailey et al. (2020) finds as income increases, people have more geographically dispersed social networks – meaning a lower proportion of their friends reside nearby. This relationship is present at all income levels: the higher someone’s income, the more dispersed their friends are. Under the assumption that social comparisons are stronger when friends are geographically closer – as their consumption decisions are likely more closely observed – this suggests that higher income individuals may experience weaker social comparisons. This results in lower centrality for higher-income communities.
Second, Chetty et al. (2022a, b) show significant homophily based on socioeconomic status, with individuals tending to befriend others with similar income levels, particularly at the upper end of the income distribution. [^17] Homophily is the tendency for people to be disproportionately connected to those similar to themselves, and is one of the most ubiquitous features of social networks (Jackson, 2008; Pin and Rogers, 2016). We then assume that people are more likely to share an identity with those of a more similar socioeconomic status. A key feature of Chetty et al.’s finding is that the level of homophily rises sharply at the very top of the income distribution – around the 90th percentile and up. This suggests that those at the very top of the income distribution may experience stronger social comparisons.
Importantly, while geographic dispersion seems to affect all income levels, the increasing homophily based on income appears primarily among those with higher incomes. So the geographic dispersion result creates a negative relationship between incomes and community density (which is driven entirely by the strengths of social comparisons) for most of the income distribution. Then the strongly increasing homophily at the top of the income distribution may create a reversal, and a positive relationship in this range. Together, these dynamics produce a non-monotonic relationship. Figure 1 provides a stylised illustration.
$w_{n}$$D_{n}$IncomeDensity
$w_{n}$$D_{n}$IncomeDensity
Figure 1: Gray line shows a possible relationship between income and density. Blue crosses (left panel) show possible realisations for Black communities. Green pluses (right panel) show possible realisations for White communities. Red arrows and red dots (both panels) show a change due to a regressive transfer between two communities.
Explaining the fact (Fact 5).
Assuming this non-monotonicity, we now consider the effect of a regressive transfer from one community to another (i.e. from a poor community to a richer one). This is a clean way of studying increases in inequality. Notice that the regressive transfer does not change the network structure – only incomes. So it will not induce any changes to density. If this regressive transfer affects two higher-income communities, then it will shift income from a lower-density community to a higher-density community. It then follows from 4(ii) that consumption in communities with unchanged incomes will decrease.
In contrast, if the regressive transfer affects two lower-income communities, then the effects will be reversed. The transfer will shift income from a higher-density to a lower-density community. This will in turn lead to an increase in consumption in unaffected communities.
If such a non-monotonic relationship exists, we would expect that Blacks are disproportionately below the trough of the density curve, while Whites are disproportionately above it, reflecting the well-documented fact that Blacks tend to have lower average incomes than Whites (Charles et al., 2009). Thus, an increase in inequality among Whites would lead to lower consumption of visible (status) goods among Whites (whose incomes did not change), while the opposite effect would be observed for Blacks. This aligns perfectly with the findings of Charles et al. (2009).
Our model can therefore explain this puzzling empirical finding without attributing it to any differences between racial groups other than well-established income disparities. we view this as a useful contribution because existing signalling models struggle to explain this empirical finding without appealing to racial differences in preferences.
6 Discrimination
As a final exercise, we extend our model to include an exogenous component to group status. To avoid confusion, we will call this component prestige. This extension explains our sixth stylised fact – that even controlling for income, different groups spend different amounts on status goods (Charles et al., 2009; Kaus, 2013). This happens because higher prestige of an agent’s group: (1) reduces her consumption of the status good, and (2) makes her consumption of the status good less responsive to her own income. So high prestige groups consume less status goods conditional on income. Given that discrimination is one determinant of prestige, our framework suggests that discrimination drives higher spending on status goods. To our knowledge, this is a novel empirical prediction.
Extending the model.
To build this into our framework, we simply adjust the definition of group status, $Y_{\theta}$. We add some exogenous prestige to the average consumption of each identity group. Formally,
$Y_{\theta}=\frac{\mathbf{x}_{\theta}+P_{\theta}}{\mathbf{x}_{-\theta}+P_{-\theta}}$
where $P_{\theta}>0$ is the prestige of group $\theta\in\{A,B\}$. Notice that group status remains relative. It now comes from two sources – consumption of the status good (an endogenous choice) and prestige (exogenous). Here, prestige is a catch-all term for everything other than consumption decisions that determines group status. [^18] Prestige itself is implicitly relative. We could capture this formally by requiring that $P_{A}+P_{B}=\text{const.}>0$. Doing so would reinforce our results – it is clear that an increase in $P_{A}$ and a decrease in $P_{B}$ push in the same direction. But it burdens the algebra without adding any insight.
The idea that there is an exogenous component to status is very much in line with the standard approach to models of social identity within economics (Akerlof and Kranton, 2000; Shayo, 2009, 2020). Implicit within this idea of an exogenous component is some underlying social structure that ascribes more prestige to some groups than to others. Our view is that one part of this social structure is discrimination – we will return to this shortly. [^19] These views, that social structures matter, and that social status has both an exogenous component and a behaviour-driven component, is also present in sociology (Bobo, 1999; Fiske, 2010; Ridgeway, 2014; Ridgeway and Markus, 2022).
Changing prestige.
We consider how a change to an identity group’s prestige affects consumption decisions by agents in both groups. The headline is straightforward: higher prestige for identity group $A$ reduces consumption by all agents in group $A$ and increases consumption by all agents in group $B$. Mechanically, this works in exactly the same way as a change in the average centrality of group $A$, except that no agent in group $A$ experiences an increase in her own centrality – which would act as a countervailing force (see 1, 2 and associated discussion). This is why the result is unambiguous for both groups.